OpenStudy (anonymous):

For what value of c is the function f(x)= {x^2-c^2 x less than or equal to 1} {cx+(5/4) 1

OpenStudy (anonymous):

This is a piecewise function

OpenStudy (anonymous):

First you need to know what continuous means. Continuous means there are no jumps or holes in the function. So this means that at the point x=1 the two parts of the piecewise function must be equal. Can you get it from here?

OpenStudy (anonymous):

I understand that, but still need some assistance in how to figure it out

OpenStudy (anonymous):

Ok just a sec and I'll give you some steps.

OpenStudy (anonymous):

Super!

OpenStudy (anonymous):

Lol I have to retype my reply since my window crashed just one more moment please.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

So you set the two functions equal to each other at x=1:\[x ^{2}-c ^{2}=cx+5/4\]\[c ^{2}+cx-x ^{2}+5/4=0\] Plugging in x=1:\[c ^{2}+c(1)+5/4-(1)^{2}=0\]\[c ^{2}+c+1/4=0\] Solve the quadratic equation for the value(s) of c that make this piecewise function continuous. Hint this is a repeated root so you only get one value. Can you take it from here?

OpenStudy (anonymous):

c=-1/2

OpenStudy (anonymous):

Thank you SO much! Stellar explanation!!!

OpenStudy (anonymous):

Yeah that's right. Good job and good luck.

OpenStudy (anonymous):

I have a question though... why do we set they functions equal to each other at x=1, if one function is x less than or equal to 1, and the other x>1?